OpenStudy (anonymous):

X-y=2 4x-3y=11 substitution method

OpenStudy (blurbendy):

x is easy to solve for here x = 2 + y now substitute 2 + y in for x in the other equation.

OpenStudy (triciaal):

so 4(2 + y) -3 y = 11 solve for y then substitute this value in x = 2 + y to get x

OpenStudy (anonymous):

Multiply out the LHS of @triciaal's equation.\[8 + 4 y - 3 y = 11 \] 4 y - 3 y = 11 - 8 y = 3 x = 2 + y x = 2 + 3 x = 5